what is the molarity and molality of a 13% solution of h2so4 solution?its density is 1.090 g/ml.to what volume should 100ml of this acid be diluted in order to prepare 1.5N solution?

Grade:12th pass

1 Answers

Naveen Kumar
9 years ago
13% solution means that 13 g of H2SO4 is in 100g of the solution.
So (100-13)g solvent contains 13 g H2SO4
.......87g...................................13gH2SO4
Hence .1g................................13/87g
...........1000g............................(13*1000/87)g H2SO4
................................................=(13*1000/98*87)mol H2SO4=molality of solution

Now density=1.090g/ml
Hence 100g solution =100/1.090 ml solution=91.74ml
So 91.74 ml of solution contains 13g i.e. 13/98 mol of H2SO4
....1ml..................................................13/(98*91.74)......
Hence, 1000ml......................................13/(98*91.74)*1000mol H2SO4=molarity of the solution=1.44M

Now basicity of this acid=2
so,normality=molarity*basicity=1.44*2=2.88N

Using N1*V1=N2*V2
2.88*100=1.5*(100+x)
x=extra volume of water used to dilute to make this solution of 1.5N